OpenStudy (anonymous):

f(u) = 2tan^-1 ( e^u) check wheather it is even function , odd or neither ?

OpenStudy (anonymous):

@shubhamsrg @sirm3d

OpenStudy (anonymous):

the function is odd

OpenStudy (shubhamsrg):

f(u) = tan-1 ( 2e^u /1-e^u) f(-u) = tan-1 (2 e^u/1-e^u) hence func is even

OpenStudy (anonymous):

@shubhamsrg the function shuld be odd...but i dont kow hw ???

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

notice that: \[\tan\theta=e^u=\frac{\sin\theta}{\cos\theta}\\ e^{-u}=\cot\theta\implies\tan^{-1}(e^{-u})=-\theta\]

OpenStudy (anonymous):

so, \[f(u)=2\tan^{-1}(e^u)=2\theta\\\implies f(-u)=2\tan^{-1}(e^{-u})=-2\theta\\\implies f(-u)=-f(u)\]

OpenStudy (shubhamsrg):

oops, I made blunder in my calculations, yes its odd.